# Estimators for multivariate allometric regression model

### v2

@@ 7 Numerical simulations
Table 1: Numerical results of the mean squared errors of $\hat{\bm{\gamma}}_{1}^{(T)}$, $\hat{\bm{\gamma}}_{1}^{(E)}$, $\hat{\bm{\gamma}}_{1}^{(R)}$, $\hat{\bm{\gamma}}_{1}(w)$ with $w=0.1,0.2,0.3,0.4,0.6$, $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ (with the average of the values of $\hat{w}^{*}$), and $\hat{\bm{\gamma}}_{1}(w^{*})$ (with the average of $w^{*}$) based on 1000 numerical simulations. The best result up to significant digits in each setting is in bold face.
(a) Results in the case under a traditional large sample regime.
$n$ $\hat{\bm{\gamma}}_{1}^{(T)}$ $\hat{\bm{\gamma}}_{1}^{(E)}$ $\hat{\bm{\gamma}}_{1}^{(R)}$ $\hat{\bm{\gamma}}_{1}(0.1)$ $\hat{\bm{\gamma}}_{1}(0.2)$ $\hat{\bm{\gamma}}_{1}(0.3)$ $\hat{\bm{\gamma}}_{1}(0.4)$ $\hat{\bm{\gamma}}_{1}(0.6)$ $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ $\hat{w}^{*}$ $\hat{\bm{\gamma}}_{1}(w^{*})$ $w^{*}$ 20 0.10517 0.90481 0.10723 0.10404 0.10113 0.09915 0.09949 0.12335 0.10101 (0.19359) 0.10109 (0.19237) 50 0.03734 0.46512 0.03887 0.03740 0.03613 0.03531 0.03540 0.04309 0.03684 (0.12519) 0.03642 (0.17417) 100 0.01831 0.22117 0.01873 0.01804 0.01745 0.01711 0.01724 0.02127 0.01782 (0.11918) 0.01762 (0.16823) 200 0.00862 0.09505 0.00886 0.00852 0.00823 0.00806 0.00812 0.01000 0.00843 (0.12590) 0.00832 (0.16535) 500 0.00349 0.03515 0.00363 0.00349 0.00337 0.00330 0.00331 0.00400 0.00344 (0.13702) 0.00341 (0.16363)
(b) Results in the case under a weak identifiability regime $\lambda_{1}-\lambda_{2}=1/n$.
$n$ $\hat{\bm{\gamma}}_{1}^{(T)}$ $\hat{\bm{\gamma}}_{1}^{(E)}$ $\hat{\bm{\gamma}}_{1}^{(R)}$ $\hat{\bm{\gamma}}_{1}(0.1)$ $\hat{\bm{\gamma}}_{1}(0.2)$ $\hat{\bm{\gamma}}_{1}(0.3)$ $\hat{\bm{\gamma}}_{1}(0.4)$ $\hat{\bm{\gamma}}_{1}(0.6)$ $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ $\hat{w}^{*}$ $\hat{\bm{\gamma}}_{1}(w^{*})$ $w^{*}$ 20 0.12627 1.44891 0.10464 0.10456 0.10508 0.10699 0.11307 0.16196 0.10721 (0.17238) 0.10461 (0.01271) 50 0.04431 1.47012 0.03730 0.03733 0.03762 0.03841 0.04022 0.05414 0.03733 (0.07300) 0.03730 (0.00440) 100 0.02267 1.46049 0.01857 0.01865 0.01888 0.01939 0.02044 0.02773 0.01860 (0.04313) 0.01857 (0.00210) 200 0.01078 1.47838 0.00905 0.00906 0.00914 0.00934 0.00979 0.01306 0.00905 (0.02688) 0.00905 (0.00102) 500 0.00442 1.48953 0.00365 0.00366 0.00371 0.00380 0.00400 0.00536 0.00365 (0.01564) 0.00365 (0.00040)
(c) Results in the case under a large $p$, large $n$ regime $n=\lfloor p^{0.8}\rfloor$, $\lambda_{1}=p^{0.8}$, $\lambda_{2}=p^{0.4}$.
$p$ $\hat{\bm{\gamma}}_{1}^{(T)}$ $\hat{\bm{\gamma}}_{1}^{(E)}$ $\hat{\bm{\gamma}}_{1}^{(R)}$ $\hat{\bm{\gamma}}_{1}(0.1)$ $\hat{\bm{\gamma}}_{1}(0.2)$ $\hat{\bm{\gamma}}_{1}(0.3)$ $\hat{\bm{\gamma}}_{1}(0.4)$ $\hat{\bm{\gamma}}_{1}(0.6)$ $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ $\hat{w}^{*}$ $\hat{\bm{\gamma}}_{1}(w^{*})$ $w^{*}$ 20 0.21682 0.63574 0.32229 0.29480 0.26881 0.24459 0.22383 0.23027 0.22046 (0.43745) 0.21935 (0.43548) 50 0.10300 0.16489 0.29513 0.21403 0.15606 0.12276 0.10712 0.10690 0.10730 (0.43157) 0.10457 (0.43582) 100 0.06570 0.08581 0.30129 0.16516 0.10299 0.07761 0.06789 0.06735 0.06790 (0.43595) 0.06662 (0.43235) 200 0.04191 0.04842 0.32524 0.11736 0.06423 0.04807 0.04296 0.04251 0.04286 (0.44012) 0.04241 (0.42782) 500 0.02412 0.02587 0.35335 0.06643 0.03371 0.02641 0.02447 0.02430 0.02451 (0.43357) 0.02433 (0.42058)
@@ Acknowledgments
This study was supported in part by Japan Society for the Promotion of Science KAKENHI Grant Numbers 20K11713 and 21K13836.

### v4

@@ 7 Numerical simulations
The best result up to significant digits in each setting is in bold face.
As for the two cases in Table 3, features of the results are different. We see from Table 3-(a) that $\hat{\bm{\gamma}}_{1}(0.3)$ is the best for $p=20$ and $\hat{\bm{\gamma}}_{1}(0.4)$ is the best for the other cases. Performances of $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ and $\hat{\bm{\gamma}}_{1}(w^{*})$ are not good but better than $\hat{\bm{\gamma}}_{1}^{(R)}$. This result may be because the upper bound in (3.1) is not sharp in high-dimensional and weakly spiked (or non-asymptotically spiked) eigenvalue situations. We see from Table 3-(b) that $\hat{\bm{\gamma}}_{1}^{(T)}$ is the best for all cases. Although $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ and $\hat{\bm{\gamma}}_{1}(w^{*})$ are not the best, their performances become closer to $\hat{\bm{\gamma}}_{1}^{(T)}$ as $p$ and $n$ increase. As for $\hat{\bm{\gamma}}_{1}^{(E)}$, its performance is bad for small $p$ and rapidly closer to $\hat{\bm{\gamma}}_{1}^{(T)}$ as $p$ and $n$ increase, which may be due to the fact from Theorem 5.1 that $\hat{\bm{\gamma}}_{1}^{(E)}$ is consistent because $\delta=0.8$ and $\beta=0.8$ satisfy $\delta>2-2\beta$. In contrast, $\hat{\bm{\gamma}}_{1}^{(R)}$ is bad especially for larger $p$, which may be due to the fact from Theorem 5.1 that $\hat{\bm{\gamma}}_{1}^{(R)}$ is not necessarily consistent because $\delta=0.8$ does not satisfy $\delta>1$.
8 Real data example
In this section, we apply the conventional methods and the proposed method to human dimensions data of Kouchi and Mochimaru (2005). In the dataset, there are 521 individuals consisting of 267 males and 254 females with 251 dimensions (variables) in common. We select $p=10$ dimensions associated with feet as response variables, and $q=5,10$ dimensions associated with hands as explanatory variables; see Table 4. The cases $q=2$-A and $q=2$-B will be explained later. In the following, we use the measurements of $n=66$ males of 19–22 years old that have no missing data on the selected dimensions.
Table 4: The dimensions selected in real data example.
(a) Response variables.
Bimalleolar breadth Medial malleolus height Sphyrion height Lateral malleolus height Sphyrion fibulare height Dorsal arch height Ball height Outside ball height Great toe tip height Foot circumference
(b) Explanatory variables. The sign $\circ$” denotes the indicated variable is used.
$q=5$ $q=10$ $q=2$-A $q=2$-B Hand length, tip to wrist crease $\circ$ $\circ$ $\circ$ Palm length $\circ$ $\circ$ Thumb length $\circ$ $\circ$ $\circ$ Index finger length $\circ$ $\circ$ $\circ$ Middle finger length $\circ$ 2nd and 3rd phalanx length, middle finger $\circ$ Middle finger length, dorsal $\circ$ Bicondylar humerus $\circ$ Bistyloid breadth $\circ$ Hand breadth, diagonal $\circ$ $\circ$ $\circ$
Remark 12.
In the case $q=10$, there is an approximate linear relationship
[…]
where the correlation coefficient between the left-hand side and the right-hand side is about 0.96. This relationship does not mess up the result because the methods to estimate $\bm{\gamma}_{1}$ discussed in this paper are based on the projections $\bm{X}(\bm{X}^{\prime}\bm{X})^{-1}\bm{X}^{\prime}\bm{Y}$ and $(\bm{I}_{n}-n^{-1}\bm{1}_{n}\bm{1}_{n}^{\prime})\bm{Y}$ of $\bm{Y}$.
Estimates of $\bm{\gamma}_{1}$ is provided in Table 5. Note that the values of $\hat{\bm{\gamma}_{1}}^{(T)}$ are the same in the two tables because explanatory variables are not used. The all coefficients in the all estimates calculated here are positive, which suggest that the first principal component represents a size factor. The contribution ratios of the first two eigenvalues are $0.393,0.264$ for $q=5$ and $0.405,0.262$ for $q=10$, so the identifiability of the first principal eigenvector does not seem weak. The values of $\hat{w}^{*}$ are $0.230$ for the case $q=5$ and $0.256$ for the case $q=10$, hence the value of $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ is different from the values of $\hat{\bm{\gamma}}_{1}^{(R)}$, $\hat{\bm{\gamma}}_{1}^{(E)}$, and $\hat{\bm{\gamma}}_{1}^{(T)}$.
Table 5: Estimates for $\gamma_{1}$ in the real data example.
(a) Results when $q=5$. The value of $\hat{w}^{*}$ is 0.230.
$\hat{\bm{\gamma}}_{1}^{(T)}$ $\hat{\bm{\gamma}}_{1}^{(E)}$ $\hat{\bm{\gamma}}_{1}^{(R)}$ $\hat{\bm{\gamma}}_{1}(0.1)$ $\hat{\bm{\gamma}}_{1}(0.2)$ $\hat{\bm{\gamma}}_{1}(0.3)$ $\hat{\bm{\gamma}}_{1}(0.4)$ $\hat{\bm{\gamma}}_{1}(0.6)$ $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ 0.181 0.159 0.188 0.187 0.186 0.184 0.183 0.178 0.185 0.329 0.371 0.319 0.321 0.322 0.324 0.326 0.333 0.323 0.305 0.371 0.288 0.290 0.293 0.296 0.300 0.312 0.294 0.245 0.209 0.247 0.248 0.247 0.247 0.247 0.244 0.247 0.165 0.044 0.206 0.201 0.194 0.186 0.176 0.151 0.191 0.202 0.336 0.150 0.157 0.166 0.176 0.188 0.218 0.169 0.076 0.101 0.065 0.066 0.068 0.070 0.073 0.079 0.069 0.034 0.028 0.036 0.036 0.035 0.035 0.035 0.033 0.035 0.058 0.072 0.052 0.053 0.054 0.055 0.056 0.059 0.054 0.792 0.725 0.803 0.802 0.801 0.799 0.796 0.787 0.800
(b) Results when $q=10$. The value of $\hat{w}^{*}$ is 0.256.
$\hat{\bm{\gamma}}_{1}^{(T)}$ $\hat{\bm{\gamma}}_{1}^{(E)}$ $\hat{\bm{\gamma}}_{1}^{(R)}$ $\hat{\bm{\gamma}}_{1}(0.1)$ $\hat{\bm{\gamma}}_{1}(0.2)$ $\hat{\bm{\gamma}}_{1}(0.3)$ $\hat{\bm{\gamma}}_{1}(0.4)$ $\hat{\bm{\gamma}}_{1}(0.6)$ $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ 0.181 0.150 0.191 0.189 0.188 0.186 0.183 0.177 0.187 0.329 0.393 0.312 0.315 0.317 0.320 0.324 0.335 0.319 0.305 0.394 0.283 0.286 0.289 0.293 0.299 0.314 0.291 0.245 0.215 0.247 0.247 0.247 0.247 0.246 0.244 0.247 0.165 0.051 0.204 0.199 0.192 0.185 0.176 0.152 0.188 0.202 0.335 0.153 0.160 0.168 0.177 0.188 0.218 0.173 0.076 0.096 0.067 0.068 0.070 0.071 0.073 0.078 0.071 0.034 0.020 0.040 0.039 0.038 0.037 0.036 0.032 0.037 0.058 0.070 0.053 0.054 0.054 0.055 0.056 0.059 0.055 0.792 0.701 0.807 0.805 0.803 0.800 0.797 0.786 0.802
In addition, we conducted the cross-validation in order to compare the proposed estimator $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ and conventional estimators from the viewpoint of prediction, which is not directly discussed in this paper. The prediction based on ordinary least squares (OLS) estimator is also considered for reference. Particularly, we calculate the values of the cross-validated mean squared prediction error
$\frac{1}{66}\sum_{i=1}^{66}\|\bm{y}_{i}-\hat{\bm{y}}_{i}\|^{2},$
where $\hat{\bm{y}}_{i}$ is a predictor for $\bm{y}_{i}$, the $i$-th row of $\bm{Y}$, constructed from the sample without $\bm{y}_{i}$ for $i=1,\ldots,66$. The result is provided in Table 6. This table shows that from the viewpoint of prediction, $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ is slightly better than the conventional estimators for the data analyzed here, while the result depends on data. Finally, two additional results with $q=2$ explanatory variables are provided; see Table 4-(b) for the dimensions selected. The cases $q=2$-A and $q=2$-B are results when good and bad sets of explanatory variables are considered, respectively. The values of $\hat{w}^{*}$ are $0.221$ and $0.468$, respectively, which implicitly suggest that the regression relatively works for the case $q=2$-A and does not work for $q=2$-B. As for the case $q=2$-A, $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ is also better than $\hat{\bm{\gamma}}_{1}^{(T)}$, $\hat{\bm{\gamma}}_{1}^{(E)}$, and $\hat{\bm{\gamma}}_{1}^{(R)}$, and OLS is the best. This result seems natural because only good explanatory variables are used, so variances becomes smaller. As for the case $q=2$-B, $\hat{\bm{\gamma}}_{1}^{(E)}$ is the best, $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ is better than $\hat{\bm{\gamma}}_{1}^{(R)}$, and $\hat{\bm{\gamma}}_{1}^{(R)}$ and OLS are bad.
Table 6: The values of the cross-validated mean squared prediction error.
$q$ 5 10 2-A 2-B
$\hat{\bm{\gamma}}_{1}^{(T)}$ 189.288 200.642 188.771 272.478
$\hat{\bm{\gamma}}_{1}^{(E)}$ 196.159 206.911 198.906 271.936
$\hat{\bm{\gamma}}_{1}^{(R)}$ 189.282 200.815 188.542 274.767
$\hat{\bm{\gamma}}_{1}(0.1)$ 189.197 200.714 188.460 273.879
$\hat{\bm{\gamma}}_{1}(0.2)$ 189.136 200.629 188.416 273.324
$\hat{\bm{\gamma}}_{1}(0.3)$ 189.112 200.573 188.428 272.949
$\hat{\bm{\gamma}}_{1}(0.4)$ 189.150 200.565 188.529 272.680
$\hat{\bm{\gamma}}_{1}(0.6)$ 189.589 200.861 189.235 272.321
$\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ 189.111 200.573 188.411 272.510
OLS 189.991 214.441 186.147 274.463
The smallest value up to significant digits in each setting is in bold face.
9 Concluding remarks
In the multivariate allometric regression model (1.2), we considered the problem of estimating $\bm{\gamma}_{1}$, a key quantity for inference of this model. Specifically, we considered a class of estimators $\hat{\bm{\gamma}}_{1}(w)$ $(0\leq w\leq 1)$ based on the weighted sum-of-squares matrices $\bm{S}_{R}$ and $\bm{S}_{E}$, established a non-asymptotic upper bound of their mean squared errors, derived the value $w^{*}$ that attains the minimum of the upper bound with respect to $w$, proposed its estimator $\hat{w}^{*}$, and discussed sufficient conditions for the consistency up to sign of $\hat{\bm{\gamma}}_{1}(w)$ $(0\leq w\leq 1)$, $\hat{\bm{\gamma}}_{1}(w^{*})$, and $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$. Based on the result of numerical simulations and a real data example, we think that $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ and $\hat{\gamma}_{1}^{(T)}=\hat{\bm{\gamma}}_{1}(0.5)$ are good from the viewpoint of estimation of $\bm{\gamma}_{1}$ and prediction when regression works. Due to a weight $w$ other than $1/2$, $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ is more flexible than $\hat{\gamma}_{1}^{(T)}$, which seems a favorable property because $w=1/2$ imposes a heavy weight on $\bm{S}_{E}$ when $\|\bm{X}\bm{\alpha}\|^{2}$ is not small and when $\lambda_{1}-\lambda_{2}$ is small, while $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ does not work well when $\lambda_{1}-\lambda_{2}$ is large but $\lambda_{1}/{\rm tr}(\bm{\Sigma})$ is small.
@@ Acknowledgments
This study was supported in part by Japan Society for the Promotion of Science KAKENHI Grant Numbers 20K11713 and 21K13836. The authors would like to thank National Institute of Advanced Industrial Science and Technology (AIST) for providing the data used in Section 8.